# GenRecal: Generation after Recalibration from Large to Small Vision-Language Models

### v2

@@ 4.2 Validation and Analysis of GenRecal
(a) Model Size and FLOPs Comparison
@@
(d) Controlling GenRecal Dataset Size
@@
(h) Compatibility with Different VLM Family
@@
(f) Comparing Cross-Tokenizer Methods
@@
(i) Compatibility with Recent VLMs
@@
With the regularization term included, as shown in Fig. 7(c), we observe that the diagonal values in the cosine similarity matrix are significantly higher than the off-diagonal values. This implies that the features of small and large VLMs are explicitly shared through Recalibrator, aligned with the results in Fig. 6 as well. In contrast, when the regularization term is removed, as depicted in Fig. 7(b), the diagonal values are comparable with the off-diagonal values. This indicates that Recalibrator fails to explicitly align large and small VLMs’ features, preventing feature matching and sharing. As a result, this leads to a significant drop in final distillation performances, as reported in Tab. 4. Insufficient alignment of large and student VLMs reduces efficient knowledge transfer during distillation. These findings highlight the crucial role of the regularization term in feature alignment and knowledge transfer for general-purpose distillation.
4.3 Training Computation and Configurations
In Tab. 5(a), we compare the model sizes and training FLOPs of the small VLM, large VLM, and Recalibrator to assess whether Recalibrator introduces additional computational overhead during training. The results show that Recalibrator’s FLOPs are substantially lower than those of both the large and small VLMs. Note that Recalibrator is removed at inference time, thus incurring no additional computational cost during inference.
We further conduct a series of ablation studies to examine various configurations affecting the distillation performance of GenRecal. As shown in Tab. 5(b), we vary the decoder depth of Recalibrator from 1 to 20 and select two depths based on a trade-off between computational efficiency and performance. Additionally, Tab. 5(c) demonstrates the importance of employing a new positional embedding (NPE) to handle newly introduced position-ids and to integrate RoPE [su2024roformer] from both large and small VLMs, which have different token types.
To analyze the effect of dataset scale, we control the dataset size from 1M to 9M samples, as shown in Tab. 5(d). The performance improvements between 5M and 9M samples are marginal, suggesting that a 5M dataset is a practical choice for users with limited training resources. Furthermore, Tab. 5(e) shows that fine-tuning the large VLM on our dataset yields further gains, indicating that the small VLM trained under the same setting has the potential to match or even surpass the large VLM’s performance.
We also compare GenRecal with recent cross-tokenizer distillation methods in Tab. 5(f). These baselines perform feature-level distillation via cross-token matching in the logit space by using classical method of Wasserstein distance [boizard2025towards] or optimal transport [cui2025multi], whereas GenRecal explicitly introduces Recalibrator to align hidden features for semantic correspondence. As observed, GenRecal achieves a substantial improvement over prior approaches.
Finally, Tab. 5(g)–(i) demonstrate GenRecal’s strong compatibility across different LLMs, VLM family, and recently released VLMs. In all cases, GenRecal consistently delivers superior performance, validating its general applicability to diverse model architectures.
@@ 4.4 Discussion
Cross-Token Matching. We can divide cross-token matching into two cases: (a) word embedding (vocabulary) matching and (b) semantic feature matching. Fundamentally, case (a) is infeasible because although the input tokens are aligned, the backbone decoder still cannot understand different word embeddings without retraining the entire model from scratch.
For case (b), UID [boizard2025towards] and MOT [cui2025multi] have recently been proposed as solutions for cross-token matching, which we compared in Tab. 5(f). However, these approaches suffer from potential information loss since they cannot properly handle token-split indices, leading to mismatched output lengths during distillation. To compensate, they just rely on zero-padding or truncation to enforce equal token lengths. In contrast, GenRecal introduces an explicit Recalibrator that produces equal-length token outputs fed into VLM-head of large VLM, thereby preventing information loss.
@@
Traditional Distillation. We further compare GenRecal with traditional distillation methods under identical token types for both teacher and student VLMs. Specifically, we use Qwen2-VL series [wang2024qwen2vl] as the teacher (72B) and student (7B or 2B) VLMs. Interestingly, even when limited to the same token types, Tab. 4 shows that GenRecal outperforms both SFT and traditional distillation by a substantial margin. Note that MiniLLM [gu2024minillm] employs reverse KL divergence, DistiLLM [ko2024distillm] uses a skewed variant, and LLaVA-KD [cai2024llava] applies normal KL divergence within a three-stage framework. These results suggest even when the large and small VLMs share identical token types, explicit feature alignment remains crucial to reduce representational disparity.
Moreover, when constructing GenRecal, VLM-head of the large VLM is exclusively used for distillation into the small VLM as illustrated in Fig. 3(a) and (b), whereas traditional distillation methods employ the student’s VLM-head to absorb the teacher’s knowledge. This design enables GenRecal to capture richer and more expressive information, as the large VLM’s VLM-head inherently possesses a higher hidden dimensionality than that of the small VLM.
@@ 5 Conclusion
Appendix A Token Type Examples
Appendix B Possible Distillation Pair for VLM Combinations: Traditional Distillation versus GenRecal
@@
Appendix C Extended Related Works
Evolution of Vision-Language Models.
Since the emergence of visual instruction tuning [liu2023visual], there surge many variations of VLMs such as LLaVA-1.5 [liu2023improved], InstructBLIP [dai2023instructblip], ShareGPT-4V [chen2023sharegpt4v], MiniGPT-v2 [chen2023minigpt] [liu2023improved, dai2023instructblip, chen2023sharegpt4v, chen2023minigpt, zhu2023minigpt], and they initially start with a standard model size of 7B. After that period, numerous VLMs [liu2024llavanext, mckinzie2024mm1, li2024mini, dong2024internlm4khd] have divided an image into multiple sub-regions to focus on the image’s details so that visual perception is enhanced. CoLLaVO [lee2024collavo] and MoAI [lee2024moai] utilize computer vision models directly for visual capability, and Mini-Gemini [li2024mini], MoVA [zong2024mova], and Eagle [shi2024eagle] employ multiple vision encoders such as CLIP [clip], ConvNext [woo2023convnext], DINO-v2 [oquab2023dinov2], and SAM [kirillov2023segment]. In parallel, Meteor [lee2024meteor] explores the efficient way of learning complex reasoning abilities, and TroL [lee2024trol] and Phantom [lee2024phantom] investigate propagation modification for how we can embed vision-language knowledge as much as possible despite using the same architectures. More recently, many VLMs like Molmo-72B [deitke2024molmo], LLaVA-OneVision-72B [li2024llava], NVLM-72B [nvlm2024], Qwen2-VL-72B [wang2024qwen2vl] and InternVL2.5-78B [chen2024expanding] have employed large-scale language models such as Qwen2/Qwen2.5-72B [yang2024qwen2]. Thanks to scaling laws [kaplan2020scaling, chung2022scaling], they have closely reached the performances of GPT-4V [gptsyscard] and Claude-3.5 Sonnet [claude3series2024]. Therefore, developing VLMs with large-scale language models is getting to a standard these days.
Knowledge Distillation.
Knowledge distillation [hinton2015distilling] provides an effective framework for transferring the rich representations of large, high-performing models to smaller, efficient counterparts while preserving performance. Early studies primarily focused on aligning the output logits of a “teacher” model with those of a “student” [sanh2019distilbert, turc2019well]. Subsequently, the FitNet framework [romero2014fitnets] extended this idea by supervising the student using intermediate feature representations rather than final outputs. In this setup, convolutional layers project the student’s features to resemble those of the teacher, and discrepancies are minimized using an L2 loss.
@@ Appendix C Extended Related Works
Appendix D Loss Graphs for Recalibrator
@@
Appendix E Implementation Details of GenRecal for Algorithms
Algorithm 1 First stage Loss Functions
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), $gt_{l}$ (label index)
[…]
7:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Algorithm 2 First stage Regularization
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), $gt_{l}$ (label index)
[…]
6:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Algorithm 3 Second stage Loss Functions
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), ($gt_{l}$, $gt_{s}$) (label index)
[…]
8:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Recalibrator’s propagation rule slightly changes depending on the hidden dimension needed. Its propagation of Fig. 3(c) and 4 correspond to that of Algorithm 1 and Algorithm 2, respectively. In Figure 3(c), Recalibrator takes small VLM’s tokens for the question and large VLM’s tokens for the answer. In contrast, in Figure 4, only large VLM’s tokens (question and answer) are fed into Recalibrator. Rec-body has the hidden dimension of the small VLM, so we should match this hidden dimension of large VLM’s tokens to that of small VLM’s ones. This is where Rec-proj-pre comes into play. This ensures compatibility with Rec-body’s hidden dimension.
Appendix F Dataset Composition and Analysis for Building GenRecal

### v3

@@ 4.3 Ablation Studies: Training Computation and Configurations
(a) Model Size and FLOPs
@@
(d) Training Dataset Size
@@
(h) Different VLM Family
@@
(f) Cross-Tokenizer Methods
@@
(i) Recent VLMs
@@
Qwen3-VL-32B [bai2025qwen3] 90.6 83.8 79.4 76.0 82.5
InternVL3.5-38B [wang2025internvl3] 90.3 81.9 82.2 76.9 82.8
@@ 4.4 Discussion
Representation Mismatch Beyond Tokenizer Mismatch. Basically, InternVL2.5 family [chen2024expanding] does not share the same tokenizer across model scales: InternVL2.5-8B uses the InternLM2.5 tokenizer [cai2024internlm2], whereas InternVL2.5-78B uses Qwen2.5 tokenizer [yang2024qwen2], as shown in Fig.˜1. On the other hands, Qwen2-VL-72B and Qwen2-VL-7B has same Qwen2 tokenizer [yang2024qwen2]. Importantly, even when the tokenizer is the same, a large parameter gap between the teacher and the student introduces a representation mismatch. Conventional distillation transfers knowledge from the large VLM head to the small VLM head, which can become a bottleneck because the student head has a much lower hidden dimensionality. In contrast, GenRecal performs distillation exclusively through the large VLM head, as illustrated in Fig.˜3. This has two advantages: (1) the large VLM head provides a higher-capacity and more expressive representation space, and (2) Recalibrator explicitly aligns the student’s hidden representations with the teacher’s feature space, serving as a representation-level warm-up before distillation. This explains why GenRecal improves performance even under the same tokenizer setting, as demonstrated in Tab.˜4. We further compare GenRecal with traditional distillation methods under identical token types for both teacher and student VLMs. Specifically, we use Qwen2-VL series [wang2024qwen2vl] as the teacher (72B) and student (7B or 2B) VLMs. Interestingly, even when limited to the same token types, Tab.˜4 shows that GenRecal outperforms both SFT and traditional distillation by a substantial margin. Note that MiniLLM [gu2024minillm] employs reverse KL divergence, DistiLLM [ko2024distillm] uses a skewed variant, and LLaVA-KD [cai2024llava] applies normal KL divergence within a three-stage framework. These results suggest even when the large and small VLMs share identical token types, explicit feature alignment remains crucial to reduce representational disparity.
Cross-Token Matching. We can divide cross-token matching into two cases: (a) word embedding (vocabulary) matching and (b) semantic feature matching. Fundamentally, case (a) is infeasible because the backbone decoder still cannot understand different word embeddings without retraining the entire model from scratch. For case (b), ULD [boizard2025towards] and MOT [cui2025multi] have recently been proposed as solutions for cross-token matching, which we compared in Tab.˜6(f). They aim to match probability mass geometry across vocabularies via token-wise Wasserstein distance or optimal transport in the logit space, but these approaches suffer from potential information loss since they cannot properly handle token-split indices, leading to mismatched output lengths during distillation. To overcome them, they just rely on zero-padding or truncation to enforce equal token lengths. In contrast, GenRecal does not perform token-level matching at the logit space. Instead, GenRecal aligns student and teacher representations before the language head through a learnable Recalibrator, enabling the teacher VLM-head to directly interpret student features. As a result, GenRecal does not require token-wise correspondence, probability sorting, padding, or transport cost design. Besides, we analyze the computation complexity and its corresponding average performance. As shown in Tab.˜6(a), Recalibrator introduces only a small fraction of FLOPs compared to both the teacher and student VLMs. Consequently, the overall FLOPs of GenRecal are comparable to recent cross-tokenizer distillation frameworks in Tab.˜6(f), but they require at least 2 epochs to reach the performance that we achieved. Based on Tab.˜7, we found that even if GenRecal requires multi training stages and new components, the additional computational overhead is comparable but achieves the highest performance. Note that we experimented them under same settings where same training dataset is used and we regard InternVL2.5-8B [chen2024expanding] as student and InternVL2.5-78B [chen2024expanding] as teacher.
Table 7: Comparison of computational cost (FLOPs and Training Time) and performance across cross-tokenizer distillation methods. We average performance across four benchmarks: MMB [liu2023mmbench], MathVista [lu2023mathvista], $M$-Vet [yu2023mm], and MMMU [yue2023mmmu].
Method Extra FLOPs Training Stages Training Time Avg. Perform.
[…]
@@ 5 Conclusion
Appendix 0.A Token Type Examples
Appendix 0.B Possible Distillation Pair for VLM Combinations: Traditional Distillation versus GenRecal
@@
Appendix 0.C Extended Related Works
Evolution of Vision-Language Models. Since the emergence of visual instruction tuning [liu2023visual], there surge many variations of VLMs such as LLaVA-1.5 [liu2023improved], InstructBLIP [dai2023instructblip], ShareGPT-4V [chen2023sharegpt4v], MiniGPT-v2 [chen2023minigpt] [liu2023improved, dai2023instructblip, chen2023sharegpt4v, chen2023minigpt, zhu2023minigpt], and they initially start with a standard model size of 7B. After that period, numerous VLMs [liu2024llavanext, mckinzie2024mm1, li2024mini, dong2024internlm4khd] have divided an image into multiple sub-regions to focus on the image’s details so that visual perception is enhanced. CoLLaVO [lee2024collavo] and MoAI [lee2024moai] utilize computer vision models directly for visual capability, and Mini-Gemini [li2024mini], MoVA [zong2024mova], and Eagle [shi2024eagle] employ multiple vision encoders such as CLIP [clip], ConvNext [woo2023convnext], DINO-v2 [oquab2023dinov2], and SAM [kirillov2023segment]. In parallel, Meteor [lee2024meteor] explores the efficient way of learning complex reasoning abilities, and TroL [lee2024trol] and Phantom [lee2024phantom] investigate propagation modification for how we can embed vision-language knowledge as much as possible despite using the same architectures. More recently, many VLMs like Molmo-72B [deitke2024molmo], LLaVA-OneVision-72B [li2024llava], NVLM-72B [nvlm2024], Qwen2-VL-72B [wang2024qwen2vl] and InternVL2.5-78B [chen2024expanding] have employed large-scale language models such as Qwen2/Qwen2.5-72B [yang2024qwen2]. Thanks to scaling laws [kaplan2020scaling, chung2022scaling], they have closely reached the performances of GPT-4V [gptsyscard] and Claude-3.5 Sonnet [claude3series2024]. Therefore, developing VLMs with large-scale language models is getting to a standard these days.
Knowledge Distillation. This provides an effective framework [hinton2015distilling] for transferring the rich representations of large, high-performing models to smaller, efficient counterparts while preserving performance. Early studies primarily focused on aligning the output logits of a “teacher” model with those of a “student” [sanh2019distilbert, turc2019well]. Subsequently, the FitNet framework [romero2014fitnets] extended this idea by supervising the student using intermediate feature representations rather than final outputs. In this setup, convolutional layers project the student’s features to resemble those of the teacher, and discrepancies are minimized using an L2 loss.
@@ Appendix 0.C Extended Related Works
Appendix 0.D Loss Graphs for Recalibrator
@@
Appendix 0.E Dataset Composition and Analysis